OpenStudy (anonymous):

What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56 ?

OpenStudy (anonymous):

\[a_1=8\] \[a_9=a_1+8d=8+8d=56\] \[8d=48\] \[d=6\] \[a_{22}=a_1+21d=8+126=134\]